OpenStudy (anonymous):

What are the x– and y–intercepts of the graph of y = x2 + x – 6?

OpenStudy (mertsj):

Find the x intercepts by letting y = 0. Find the y intercept by letting x = 0

OpenStudy (anonymous):

x^2+x-6 Delta = 1-4(1)(-6)=25 x1=-6/2=-3 x2=-4/2=-2 the x's intersepts 0^2+0-6=-6 Y intersept

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Your welcome